OpenStudy (anonymous):

Solve following the System of Equation: -4x + y = -1 2x + y = 5

OpenStudy (shamil98):

-4x + y = -1 4x + 2y = 10 (you multiply the second equation by 2) from there add the equations and solve for y

OpenStudy (unklerhaukus):

after you get y, (following shamil98s method) substitute y into one of the original equation to solve for x